A 12m lever is used to lift 1000N load fixed at one end 2m from the fulcrum.what is the effort required to lift the load exerted at the other end.assume efficiency is 100%


1
Expert's answer
2021-07-22T13:06:05-0400

The equation for lever arm is the following (see https://www.youtube.com/watch?v=jg4e8W44_E4):


"F_1l_1 = F_2l_2"

where "F_1 = 1000N" is the load fixed at "l_1 = 2m" from the fulcrum, and "F_2" is the effort applied to another end of the lever at "l_2 = 12m-2m = 10m" from the fulcrum. Thus, obtain:


"F_2 = \dfrac{F_1l_1}{l_2} = \dfrac{1000N\cdot 2m}{10m} = 200N"

Answer. 200N.